OpenStudy (lgbasallote):

how many words can be formed from LGBASALLOTE? words not necessarily to be in the dictionary.. i dont suppose this is a simple permutation problem? and yeah im narcissistic so i always use my name :p

OpenStudy (anonymous):

Letters: L 3 of 'em A 2 G 1 B 1 S 1 O 1 T 1 E 1 So, 11 total. Counting 11 character words ain't so bad. Let's place the Ls and the As first. Number of ways to place the Ls: 11 choose 3 Now that I've chosen the Ls, the number of ways to place the As: 8 choose 2 Now that I've chosen the Ls and As, I need only choose a letter for the remaining 6 spots which is 6! (11C3)*(8C2)*(6!)

OpenStudy (lgbasallote):

the C means...?

OpenStudy (anonymous):

"Choose" Combinations. \[(11C3) = \frac{11!}{(11-3)!3!}\]

OpenStudy (lgbasallote):

so how do i interpret the expression you had in the \(\frac{n!}{r!}\) form?

OpenStudy (anonymous):

You want this expanded? (11C3)*(8C2)*(6!)

OpenStudy (anonymous):

\[\frac{11!}{(11-3)!3!}\frac{8!}{(8-2)!2!}6!\]

OpenStudy (lgbasallote):

wait wait wait...isnt \[\large \frac{11!}{8!3!} = 1??\]

OpenStudy (lgbasallote):

oh wait no sorry

OpenStudy (lgbasallote):

but let's say i choose A first will it be the same as \[\Large \frac{11!}{9!2!} \times \frac{8!}{5!3!} \times 6!\]

OpenStudy (lgbasallote):

the second numerator is supposed to be 9!

OpenStudy (anonymous):

Right. And 6!3! in the denominator

OpenStudy (lgbasallote):

right..exactly what i meant haha thanks

OpenStudy (anonymous):

Cool, so yeah. That gives us all of the possible 11 character long words... it doesn't tell us all of the possibilities for shorter words, which becomes a pretty complex problem. Doable, but a bit tedious.

OpenStudy (anonymous):

11!/2!3! ?

OpenStudy (lgbasallote):

wait...the possible 11 character words...isnt it supposed to be \[\large \frac{11!}{3!2!}\]?

OpenStudy (lgbasallote):

oh lol it is...

OpenStudy (lgbasallote):

i thought that was the possible number of words...that was just an expanded form of the number of how many times my name can be arranged :p

OpenStudy (anonymous):

Hmm okay. Yeah, that's a somewhat more elegant way to calculate it.

OpenStudy (anonymous):

"LGBASALLOTE" has 11 letters among which 3 L's and 2 A's the rest are all different. Total number of permutation possible taken all at a time: \[ \frac{11!}{3!\times 2!}\]

OpenStudy (lgbasallote):

well i knew about the part of the 11 characters...no way to tell for the smaller characters?

OpenStudy (anonymous):

Only 10 Characters: Option 1: only 2 Ls- \[\frac{10!}{2!2!}\] Option 2: only 1 A- \[\frac{10!}{3!}\] Option 3: 3Ls 2As, 1 of each other letter besides 1 \[\frac{10!}{3!2!}*6\] Add those up to get the number of 10 char words.

OpenStudy (lgbasallote):

ugh o.O that's enough to show me it wont be simple :p lol

OpenStudy (anonymous):

Do you follow the logic there?

OpenStudy (lgbasallote):

somehow...

OpenStudy (anonymous):

Haha okay good.